Leann just sold a $10,000 par value bond for $9,800. The bond interest rate was 5.5% per year payable quarterly. Leann owned the bond for 3 years. The 1st interest payment she received was 3 months after she bought the bond. She sold it immediately after receiving her 12th interest payment. Leann’s yield on the bond was 10.5% per year compounded quarterly. Determine the price she paid when she purchased the bond.
Answer #1

quarterly Interest = 5.5%*10000/4 = 137.50
Rate per quarter = 10.5%/4
Price of Bond = PV of Coupons + PV of Sales Price = 137.50*(1-(1+10.5%/4)-12)/10.5% + 9800/(1+10.5%/4)12 = 8580.88

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well